# Towards Foundation Model for Chemical Reactor Modeling: Meta-Learning with Physics-Informed Adaptation

### v2

@@ 4.1 CSTR Modeling and Physics-Informed Loss Term
$\frac{dC_{A}}{dt}=$ $\frac{F}{V}(C_{A0}-C_{A})-k_{0}e^{\frac{-E}{RT}}C_{A}^{2}$ (4a)
$\frac{dT}{dt}=$ $\frac{F}{V}(T_{0}-T)+\frac{-\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E}{RT}}C_{A}^{2}+\frac{Q}{\rho_{L}C_{p}V}$ (4b)
where $C_{A}$ is the concentration of reactant A, $V$ is the volume of the reacting liquid, $T$ is the temperature, $Q$ is the heat input rate, $C_{A0}$ is the inlet concentration of reactant A, $T_{0}$ is the inlet temperature, $F$ is the volumetric flow rate, $\rho_{L}$ is the constant density of the reacting liquid, $C_{p}$ is the heat capacity, $\Delta H$ is the enthalpy of reaction, $k_{0}$ is the pre-exponential constant, $E$ is the activation energy, and $R$ is the ideal gas constant. The manipulated inputs (i.e., control) in this system are represented by $\Delta C_{A0}=C_{A0}-C_{A0_{s}}$ and $\Delta Q=Q-Q_{s}$, such that the equilibrium point of the system is located at the origin of the state-space. The ranges of parameter values for the CSTRs are described in Table 1.
@@
In the context of CSTR modeling, the inputs consist of $[T_{t},C_{A_{t}},\Delta Q_{t},\Delta C_{{A0}_{t}}]$ at the current time step $t$. The outputs entail the state trajectory $[T_{t+1,\ldots,n},C_{A_{t+1,\ldots,n}}]$ over the subsequent $n$ time steps. Leveraging randomly drawn parameter values from a reasonable range and employing the material and energy balance equations depicted in Eq. (4), we perform numerical simulations of various CSTRs to collect their respective output sequences employing the explicit Euler method. The operational ranges of CSTRs and the parameters used for simulation are as follows: $T$, $C_{A}$, $\Delta Q$, $\Delta C_{A0}$ are evenly sampled from $[300,600]~{}K$, $[0,6]~{}kmol/m^{3}$, $[-5,5]\times 10^{5}~{}kJ/hr$, and $[-3.5,3.5]~{}kmol/m^{3}$ respectively. For the explicit Euler method, the integration time step $h_{c}=1\times 10^{-4}~{}hr$ is used to solve Eq. (4), and the manipulated inputs $\Delta Q$, $\Delta C_{A0}$ are changed every sampling period of $5\times 10^{-3}~{}hr$. However, it is noted that the ranges of variables can be chosen appropriately for the foundation model depending on the operating conditions of interest.
Subsequently, when implementing the foundation model to a new process (e.g., a new reaction in a CSTR), we will apply physics-informed adaptation method. Specifically, the physics-informed loss function for learning a new CSTR is the weighted sum of the data-driven loss term $\mathcal{L}_{d}$ and the physics-informed loss term $\mathcal{L}_{p}$. $\mathcal{L}_{d}$ takes the form of MSE of the actual output $y$ and the predicted output $\tilde{y}$ using the actual few-shot data, as follows:
@@
$\mathcal{L}_{C_{A_{t+i}}}=$ $\frac{d\widetilde{C_{A}}_{t+i}}{dt}-\frac{F}{V}(C_{A0_{t}}-\widetilde{C_{A}}_{t+i})+k_{0}e^{\frac{-E}{R\widetilde{T}_{t+i}}}\widetilde{C_{A}}_{t+i}^{2}$ (7a)
$\mathcal{L}_{T_{t+i}}=$ $\frac{d\widetilde{T}_{t+i}}{dt}-\frac{F}{V}(T_{0}-\widetilde{T}_{t+i})+\frac{\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E}{R\widetilde{T}_{t+i}}}\widetilde{C_{A}}_{t+i}^{2}-\frac{Q_{t}}{\rho_{L}C_{p}V}$ (7b)
Here, $\widetilde{C_{A}}$ and $\widetilde{T}$ represent the predicted outputs corresponding to the collocation points as inputs and we use the explicit Euler method to solve the ODEs. When operating a real-world CSTR, certain parameters such as $F$, $V$, $T_{0}$, $C_{A0_{s}}$, and $Q_{s}$ are typically known through expert knowledge and direct physical measurements. However, other critical parameters, such as $k_{0}$, $E$, $\Delta H$, $C_{p}$, and $\rho_{L}$, may not be directly measurable and therefore remain unknown. To compute $\mathcal{L}_{p}$, we utilize the actual values of $F$, $V$, $T_{0}$, $C_{A0_{s}}$, and $Q_{s}$, while estimating the values of $k_{0}$, $E$, $\Delta H$, $C_{p}$, and $\rho_{L}$ in Eq. (7). Specifically, we use the median values of $k_{0}$, $E$, $\Delta H$, $C_{p}$, and $\rho_{L}$ from their respective sampling ranges.
@@ 4.2 BR Modeling and Physics-Informed Loss Term
$\frac{dC_{A}}{dt}=$ $-k_{0}e^{\frac{-E}{RT}}C_{A}$ (8a)
$\frac{dT}{dt}=$ $\frac{-\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E}{RT}}C_{A}+\frac{Q}{\rho_{L}C_{p}V}$ (8b)
@@
In the context of BR modeling, to maintain uniform input dimensions akin to CSTRs, we implement padding on the inputs for BRs. Consequently, the final inputs comprise $[T_{t},C_{A_{t}},\Delta Q_{t},\Delta Q_{t}]$ at the current time step $t$, while the outputs encapsulate the state trajectory $[T_{t+1,\ldots,n},C_{A_{t+1,\ldots,n}}]$ over the subsequent $n$ time steps. Leveraging randomly drawn parameter values from a reasonable range and employing the material and energy balance equations depicted in Eq. (8), we conduct numerical simulations of various BRs to collect their respective output sequences employing the explicit Euler method. Specifically, the operational ranges of BRs as well as the parameters used for data simulation are as follows: $T$, $C_{A}$, $\Delta Q$ are evenly sampled from $[300,600]~{}K$, $[0,6]~{}kmol/m^{3}$, and $[-5,5]\times 10^{5}~{}kJ/hr$, respectively. For explicit Euler method, the integration time step $h_{c}=1\times 10^{-4}~{}hr$ is used to solve Eq. (8), and the manipulated input $\Delta Q$ is changed every sampling period of $5\times 10^{-2}~{}hr$.
@@
$\mathcal{L}_{C_{A_{t+i}}}=$ $\frac{d\widetilde{C_{A}}_{t+i}}{dt}+k_{0}e^{\frac{-E}{R\widetilde{T}_{t+i}}}\widetilde{C_{A}}_{t+i}$ (9a)
$\mathcal{L}_{T_{t+i}}=$ $\frac{d\widetilde{T}_{t+i}}{dt}+\frac{\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E}{R\widetilde{T}_{t+i}}}\widetilde{C_{A}}_{t+i}-\frac{Q_{t}}{\rho_{L}C_{p}V}$ (9b)
Given a new BR, we also assume that the actual values of $V$ and $Q$ are known, while the values of $k_{0}$, $E$, $\Delta H$, $C_{p}$, and $\rho_{L}$ remain unknown and will be estimated (i.e., using the median values of their respective sampling ranges) in Eq. (9) for model adaptation.
@@ 4.3 PFR Modeling and Physics-Informed Loss Term
$\frac{\partial C_{A}}{\partial t}=$ $-u\frac{\partial C_{A}}{\partial z}-k_{0}e^{\frac{-E}{RT}}C_{A}$ (10a)
$\frac{\partial T}{\partial t}=$ $-u\frac{\partial T}{\partial z}+\frac{-\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E}{RT}}C_{A}+\frac{U}{\rho_{L}C_{p}A}A_{c}(T_{c}-T)$ (10b)
@@
$\frac{dC_{A_{N}}}{dt}=$ $-u\frac{C_{A_{N}}-C_{A_{N-1}}}{Z_{N-1}-Z_{N}}-k_{0}e^{\frac{-E}{RT}}C_{A}$ (11a)
$\frac{dT_{N}}{dt}=$ $-u\frac{T_{N}-T_{N-1}}{Z_{N-1}-Z_{N}}+\frac{-\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E}{RT}}C_{A}+\frac{U}{\rho_{L}C_{p}A}A_{c}(T_{c}-T)$ (11b)
@@
In the context of PFR modeling, to maintain uniform input dimensions akin to CSTRs and BRs, we implement padding on the inputs for PFRs. Furthermore, we exclusively monitor the output from a single discretized location over the sampling period. Specifically, in this study, we focus on tracking the output from the first discretized point, where $N=1$. Consequently, the final inputs comprise $[T_{t},C_{A_{t}},\Delta T_{c_{t}},\Delta T_{c_{t}}]$ at the current time step $t$, while the outputs encapsulate the state trajectory $[T_{t+1,\ldots,n},C_{A_{t+1,\ldots,n}}]$ over the subsequent $n$ time steps. Leveraging parameter values randomly drawn from a reasonable range and employing the material and energy balance equations depicted in Eq. (11), we perform numerical simulations of various PFRs to collect their respective output sequences, which involve employing the explicit Euler method. Specifically, the operational ranges of PFRs as well as the parameters used for data simulation are as follows: $T$, $C_{A}$, $\Delta T_{c}$ are evenly sampled from $[300,500]~{}K$, $[0.5,3]~{}kmol/m^{3}$, and $[100,300]~{}K$, respectively. For explicit Euler method, the integration time step $h_{c}=0.01~{}hr$ is used to solve Eq. (11), and the manipulated input $\Delta T_{c}$ is changed every sampling period of $0.1~{}hr$.
@@
$\mathcal{L}_{C_{A_{t+i}}}=$ $\frac{d\widetilde{C_{A}}_{t+i}}{dt}+u\frac{\widetilde{C_{A}}_{t+i}-C_{A_{t}}}{Z_{1}-Z_{0}}+k_{0}e^{\frac{-E}{R\widetilde{T}_{t+i}}}\widetilde{C_{A}}_{t+i}$ (12a)
$\mathcal{L}_{T_{t+i}}=$ $\frac{d\widetilde{T}_{t+i}}{dt}+u\frac{\widetilde{T}_{t+i}-T_{t}}{Z_{1}-Z_{0}}-\frac{-\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E}{R\widetilde{T}_{t+i}}}\widetilde{C_{A}}_{t+i}-\frac{U}{\rho_{L}C_{p}A}A_{c}(T_{c_{t}}-\widetilde{T}_{t+i})$ (12b)
The calculation of the physics-informed loss term follows the method for CSTRs and BRs. Specifically, we utilize the actual values of $L$, $N$, $T_{c_{s}}$, $A$, $A_{c}$, $u$, and $U$, and estimate the values of $k_{0}$, $E$, $\Delta H$, $C_{p}$, and $\rho_{L}$ (i.e., using the median values of $k_{0}$, $E$, $\Delta H$, $C_{p}$, and $\rho_{L}$ from their respective sampling ranges) in Eq. (12) for model adaptation. Note that the terms $\frac{d\widetilde{C_{A}}_{t+i}}{dt}$ and $\frac{d\widetilde{T}_{t+i}}{dt}$ in $\mathcal{L}_{p}$ of CSTRs, BRs, and PFRs, can be computed using numerical methods such as the explicit Euler method.
@@
We apply padding to the inputs of BRs and PFRs to ensure a consistent input dimension for the foundation model, accommodating all three types of reactors. Ablation studies revealed that padding with manipulated inputs (e.g., $\Delta Q$ for BRs and $\Delta T_{c}$ for PFRs) improves model performance compared to conventional padding methods, such as using natural values, e.g., $-1$, 00, or $1$.
@@
Training PINNs can be notably challenging, and their performance is highly sensitive to the hyperparameter $\gamma$ governing the weighting of the loss function. In summary, the loss term of CSTR, BR, and PFR can be expressed as follows:
@@
where $\gamma_{1}=1\times 10^{3}$, $\gamma_{2}=1\times 10^{-2}$, and $\gamma_{3}=1\times 10^{-5}$ give the best performance for all three types of reactors in our case.
@@
It is essential to highlight that we enforce constraints on the simulated output trajectories for all CSTRs, BRs, and PFRs to ensure that they are within a plausible range. This measure is crucial to ensure that the simulated CSTRs, BRs, and PFRs adhere to theoretical validity. Furthermore, the integration time step and the sampling period vary across CSTRs, BRs, and PFRs in this simulation study due to different kinetics and process dynamics. However, it is important to note that the foundation model can accommodate differing integration time steps and sampling periods for real-world reactors, provided the input dimensions remain consistent. This is due to the fact that the inputs to the neural network are merely numerical values, lacking any inherent physical interpretation for the network.

### v3

@@ 4 Data Simulation for Chemical Reactors and Development of Physics-Informed Loss Term
Regarding the chemistry of the system, our study does not focus on a specific chemical reaction but rather on a general class of reactions where species $A$ is converted into species $B$ (e.g., dimerization of Nitrogen Dioxide to Dinitrogen Tetroxide, dimerization of $\alpha$-Methylstyrene, dimerization of Isoprene, dimerization of Formaldehyde to Paraformaldehyde, etc.). This benchmark reaction is widely used in process modeling due to its nonlinear dynamics and the challenges associated with accurately characterizing reaction kinetics, particularly when precise kinetic parameters are unavailable. Our “foundation model” is designed to learn the behavior of a broad range of elementary reactions (integer-order) with kinetic parameters sampled within reasonable ranges. The objective is not to develop a model for a single, well-defined reaction but rather to create a generalizable framework adaptable to various chemical systems.
@@ 4.1 CSTR Modeling and Physics-Informed Loss Term
$\frac{dC_{A}}{dt}=$ $\frac{F}{V}(C_{A0}-C_{A})-k_{0}e^{\frac{-E_{a}}{RT}}C_{A}^{2}$ (4a)
$\frac{dT}{dt}=$ $\frac{F}{V}(T_{0}-T)+\frac{-\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E_{a}}{RT}}C_{A}^{2}+\frac{Q}{\rho_{L}C_{p}V}$ (4b)
where $C_{A}$ is the concentration of reactant A, $V$ is the volume of the reacting liquid, $T$ is the temperature, $Q$ is the heat input rate, $C_{A0}$ is the inlet concentration of reactant A, $T_{0}$ is the inlet temperature, $F$ is the volumetric flow rate, $\rho_{L}$ is the constant density of the reacting liquid, $C_{p}$ is the heat capacity, $\Delta H$ is the enthalpy of reaction, $k_{0}$ is the pre-exponential constant, $E_{a}$ is the activation energy, and $R$ is the ideal gas constant. The manipulated inputs (i.e., control) in this system are represented by $\Delta C_{A0}=C_{A0}-C_{A0_{s}}$ and $\Delta Q=Q-Q_{s}$, such that the equilibrium point of the system is located at the origin of the state-space. The ranges of parameter values for the CSTRs are described in Table 1.
Remark 7.
In a chemical reactor (e.g., CSTR, PFR), an equilibrium point typically refers to a steady-state operating condition, where key state variables (e.g., concentrations, temperature) remain constant over time. In an ODE-based reactor model,
$\frac{dx}{dt}=f(x,u)$ (5)
where $x$ represents state variables (e.g., concentration, temperature) and $u$ represents inputs (e.g., flow rates, heating), the equilibrium point $x^{*}$ satisfies $f(x^{*},u^{*})=0$, meaning the time derivatives vanish. This equilibrium point can be determined by solving the steady-state form of the reactor model (i.e., setting the governing equations to zero) or by numerically identifying fixed points of the system. A more detailed analysis can be found in our previous work (Wu et al., 2019a).
@@
In the context of CSTR modeling, the inputs consist of $[T_{t},C_{A_{t}},\Delta Q_{t},\Delta C_{{A0}_{t}}]$ at the current time step $t$. The outputs entail the state trajectory $[T_{t+1,\ldots,n},C_{A_{t+1,\ldots,n}}]$ over the subsequent $n$ time steps. Leveraging randomly drawn parameter values from a reasonable range and employing the material and energy balance equations depicted in Eq. (4), we perform numerical simulations of various CSTRs to collect their respective output sequences employing the explicit Euler method. The operational ranges of CSTRs and the parameters used for simulation are as follows: $T$, $C_{A}$, $\Delta Q$, $\Delta C_{A0}$ are evenly sampled from $[300,600]~{}K$, $[0,6]~{}kmol/m^{3}$, $[-5,5]\times 10^{5}~{}kJ/hr$, and $[-3.5,3.5]~{}kmol/m^{3}$ respectively. For the explicit Euler method, the integration time step $h_{c}=1\times 10^{-4}~{}hr$ is used to solve Eq. (4), and the manipulated inputs $\Delta Q$, $\Delta C_{A0}$ are changed every sampling period of $5\times 10^{-3}~{}hr$. However, it is noted that the ranges of variables can be chosen appropriately for the “foundation model” depending on the operating conditions of interest.
Remark 8.
We have tested different numerical integration methods, including Runge-Kutta and implicit Euler, and found that the differences in results are minimal. This is particularly true for the CSTR system, which we previously analyzed and confirmed is not significantly stiff (Wu et al., 2019a). Additionally, we ensured that the chosen time step is sufficiently small, preventing numerical oscillations or divergence. Thus, the explicit Euler method provides reliable training data for our case.
Subsequently, when implementing the “foundation model” to a new process (e.g., a new reaction in a CSTR), we will apply physics-informed adaptation method. Specifically, the physics-informed loss function for learning a new CSTR is the weighted sum of the data-driven loss term $\mathcal{L}_{d}$ and the physics-informed loss term $\mathcal{L}_{p}$. $\mathcal{L}_{d}$ takes the form of MSE of the actual output $y$ and the predicted output $\tilde{y}$ using the actual few-shot data, as follows:
@@
$\mathcal{L}_{C_{A_{t+i}}}=$ $\frac{d\widetilde{C_{A}}_{t+i}}{dt}-\frac{F}{V}(C_{A0_{t}}-\widetilde{C_{A}}_{t+i})+k_{0}e^{\frac{-E_{a}}{R\widetilde{T}_{t+i}}}\widetilde{C_{A}}_{t+i}^{2}$ (8a)
$\mathcal{L}_{T_{t+i}}=$ $\frac{d\widetilde{T}_{t+i}}{dt}-\frac{F}{V}(T_{0}-\widetilde{T}_{t+i})+\frac{\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E_{a}}{R\widetilde{T}_{t+i}}}\widetilde{C_{A}}_{t+i}^{2}-\frac{Q_{t}}{\rho_{L}C_{p}V}$ (8b)
Here, $\widetilde{C_{A}}$ and $\widetilde{T}$ represent the predicted outputs corresponding to the collocation points as inputs and we use the explicit Euler method to solve the ODEs. When operating a real-world CSTR, certain parameters such as $F$, $V$, $T_{0}$, $C_{A0_{s}}$, and $Q_{s}$ are typically known through expert knowledge and direct physical measurements. However, other critical parameters, such as $k_{0}$, $E_{a}$, $\Delta H$, $C_{p}$, and $\rho_{L}$, may not be directly measurable and therefore remain unknown. To compute $\mathcal{L}_{p}$, we utilize the actual values of $F$, $V$, $T_{0}$, $C_{A0_{s}}$, and $Q_{s}$, while estimating the values of $k_{0}$, $E_{a}$, $\Delta H$, $C_{p}$, and $\rho_{L}$ in Eq. (8). Specifically, we use the median values of $k_{0}$, $E_{a}$, $\Delta H$, $C_{p}$, and $\rho_{L}$ from their respective sampling ranges.
@@ 4.2 BR Modeling and Physics-Informed Loss Term
$\frac{dC_{A}}{dt}=$ $-k_{0}e^{\frac{-E_{a}}{RT}}C_{A}$ (9a)
$\frac{dT}{dt}=$ $\frac{-\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E_{a}}{RT}}C_{A}+\frac{Q}{\rho_{L}C_{p}V}$ (9b)
@@
In the context of BR modeling, to maintain uniform input dimensions akin to CSTRs, we implement padding on the inputs for BRs. Consequently, the final inputs comprise $[T_{t},C_{A_{t}},\Delta Q_{t},\Delta Q_{t}]$ [^1] It is important to highlight that the variable $\Delta Q_{t}$ is not repeated; rather, it appears due to padding along the feature dimension to ensure uniform input size for the model. This padding is necessary for processing sequences in the RNN while maintaining a consistent feature representation across different samples. at the current time step $t$, while the outputs encapsulate the state trajectory $[T_{t+1,\ldots,n},C_{A_{t+1,\ldots,n}}]$ over the subsequent $n$ time steps. Leveraging randomly drawn parameter values from a reasonable range and employing the material and energy balance equations depicted in Eq. (9), we conduct numerical simulations of various BRs to collect their respective output sequences employing the explicit Euler method. Specifically, the operational ranges of BRs as well as the parameters used for data simulation are as follows: $T$, $C_{A}$, $\Delta Q$ are evenly sampled from $[300,600]~{}K$, $[0,6]~{}kmol/m^{3}$, and $[-5,5]\times 10^{5}~{}kJ/hr$, respectively. For explicit Euler method, the integration time step $h_{c}=1\times 10^{-4}~{}hr$ is used to solve Eq. (9), and the manipulated input $\Delta Q$ is changed every sampling period of $5\times 10^{-2}~{}hr$.
@@
$\mathcal{L}_{C_{A_{t+i}}}=$ $\frac{d\widetilde{C_{A}}_{t+i}}{dt}+k_{0}e^{\frac{-E_{a}}{R\widetilde{T}_{t+i}}}\widetilde{C_{A}}_{t+i}$ (10a)
$\mathcal{L}_{T_{t+i}}=$ $\frac{d\widetilde{T}_{t+i}}{dt}+\frac{\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E_{a}}{R\widetilde{T}_{t+i}}}\widetilde{C_{A}}_{t+i}-\frac{Q_{t}}{\rho_{L}C_{p}V}$ (10b)
Given a new BR, we also assume that the actual values of $V$ and $Q$ are known, while the values of $k_{0}$, $E_{a}$, $\Delta H$, $C_{p}$, and $\rho_{L}$ remain unknown and will be estimated (i.e., using the median values of their respective sampling ranges) in Eq. (10) for model adaptation.
@@ 4.3 PFR Modeling and Physics-Informed Loss Term
$\frac{\partial C_{A}}{\partial t}=$ $-u\frac{\partial C_{A}}{\partial z}-k_{0}e^{\frac{-E_{a}}{RT}}C_{A}$ (11a)
$\frac{\partial T}{\partial t}=$ $-u\frac{\partial T}{\partial z}+\frac{-\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E_{a}}{RT}}C_{A}+\frac{U}{\rho_{L}C_{p}A}A_{c}(T_{c}-T)$ (11b)
@@
$\frac{dC_{A_{N}}}{dt}=$ $-u\frac{C_{A_{N}}-C_{A_{N-1}}}{Z_{N-1}-Z_{N}}-k_{0}e^{\frac{-E_{a}}{RT}}C_{A}$ (12a)
$\frac{dT_{N}}{dt}=$ $-u\frac{T_{N}-T_{N-1}}{Z_{N-1}-Z_{N}}+\frac{-\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E_{a}}{RT}}C_{A}+\frac{U}{\rho_{L}C_{p}A}A_{c}(T_{c}-T)$ (12b)
@@
In the context of PFR modeling, to maintain uniform input dimensions akin to CSTRs and BRs, we implement padding on the inputs for PFRs. Furthermore, we exclusively monitor the output from a single discretized location over the sampling period. Specifically, in this study, we focus on tracking the output from the first discretized point, where $N=1$. Consequently, the final inputs comprise $[T_{t},C_{A_{t}},\Delta T_{c_{t}},\Delta T_{c_{t}}]$ [^2] It is important to highlight that the variable $\Delta T_{c_{t}}$ is not repeated; rather, it appears due to padding along the feature dimension to ensure uniform input size for the model. at the current time step $t$, while the outputs encapsulate the state trajectory $[T_{t+1,\ldots,n},C_{A_{t+1,\ldots,n}}]$ over the subsequent $n$ time steps. Leveraging parameter values randomly drawn from a reasonable range and employing the material and energy balance equations depicted in Eq. (12), we perform numerical simulations of various PFRs to collect their respective output sequences, which involve employing the explicit Euler method. Specifically, the operational ranges of PFRs as well as the parameters used for data simulation are as follows: $T$, $C_{A}$, $\Delta T_{c}$ are evenly sampled from $[300,500]~{}K$, $[0.5,3]~{}kmol/m^{3}$, and $[100,300]~{}K$, respectively. For explicit Euler method, the integration time step $h_{c}=0.01~{}hr$ is used to solve Eq. (12), and the manipulated input $\Delta T_{c}$ is changed every sampling period of $0.1~{}hr$.
@@
$\mathcal{L}_{C_{A_{t+i}}}=$ $\frac{d\widetilde{C_{A}}_{t+i}}{dt}+u\frac{\widetilde{C_{A}}_{t+i}-C_{A_{t}}}{Z_{1}-Z_{0}}+k_{0}e^{\frac{-E_{a}}{R\widetilde{T}_{t+i}}}\widetilde{C_{A}}_{t+i}$ (13a)
$\mathcal{L}_{T_{t+i}}=$ $\frac{d\widetilde{T}_{t+i}}{dt}+u\frac{\widetilde{T}_{t+i}-T_{t}}{Z_{1}-Z_{0}}-\frac{-\Delta H}{\rho_{L}C_{p}}k_{0}e^{\frac{-E_{a}}{R\widetilde{T}_{t+i}}}\widetilde{C_{A}}_{t+i}-\frac{U}{\rho_{L}C_{p}A}A_{c}(T_{c_{t}}-\widetilde{T}_{t+i})$ (13b)
The calculation of the physics-informed loss term follows the method for CSTRs and BRs. Specifically, we utilize the actual values of $L$, $N$, $T_{c_{s}}$, $A$, $A_{c}$, $u$, and $U$, and estimate the values of $k_{0}$, $E_{a}$, $\Delta H$, $C_{p}$, and $\rho_{L}$ (i.e., using the median values of $k_{0}$, $E_{a}$, $\Delta H$, $C_{p}$, and $\rho_{L}$ from their respective sampling ranges) in Eq. (13) for model adaptation. Note that the terms $\frac{d\widetilde{C_{A}}_{t+i}}{dt}$ and $\frac{d\widetilde{T}_{t+i}}{dt}$ in $\mathcal{L}_{p}$ of CSTRs, BRs, and PFRs, can be computed using numerical methods such as the explicit Euler method.
@@
We apply padding to the inputs of BRs and PFRs to ensure a consistent input dimension for the “foundation model”, accommodating all three types of reactors. Here, padding is not intended to improve results but rather to ensure uniform input dimensions, which is crucial for processing data in an RNN. Specifically, we pad along the feature dimension to maintain consistency across inputs. Without this, we would need to train separate RNNs for different input types, which contradicts our goal of developing a single unified model capable of handling various chemical reactions across different reactor types. Ablation studies revealed that padding with manipulated inputs (e.g., $\Delta Q$ for BRs and $\Delta T_{c}$ for PFRs) improves model performance compared to conventional padding methods, such as using natural values, e.g., $-1$, 00, or $1$.
@@
Training PINNs can be notably challenging, and their performance is highly sensitive to the hyperparameters $\gamma_{i}$ governing the weighting of the loss terms. The weighting of the loss components was chosen to ensure that all loss terms remain within the same order of magnitude - a common practice in multi-objective optimization. Specifically, we assign a weight of $\gamma_{1}=1\times 10^{3}$ to the data-driven loss $\mathcal{L}_{d}$, reflecting our greater reliance on it as the primary learning signal. In contrast, the physics-informed losses $\mathcal{L}_{C_{A}}$ and $\mathcal{L}_{T}$ are assigned smaller weights, $\gamma_{2}=1\times 10^{-2}$, and $\gamma_{3}=1\times 10^{-5}$, respectively. This design reflects our intention for the model to primarily learn from data, with physics-based constraints serving as informative regularization rather than dominating the training process. These weights were selected based on empirical tuning through grid search over a predefined range (i.e., $\gamma_{i}\in[1\times 10^{-5},1\times 10^{5}])$, selecting the combination that yielded the lowest validation error. In summary, the total loss for CSTR, BR, and PFR models is formulated as:
@@
The chosen weights gave the best performance across the three reactor types within the explored hyperparameter space, although they may not represent a global optimum.
@@
It is essential to highlight that we enforce constraints on the simulated output trajectories for all CSTRs, BRs, and PFRs to ensure that they remain within a physically plausible range. Specifically, we applied checks during data generation to detect and remove trajectories containing unrealistic values such as Inf (infinity), NaN (not-a-number), negative concentrations or temperatures (where physically infeasible), extremely large values, or values close to zero where nonzero quantities are expected. If a generated trajectory violated any of these conditions, we excluded the corresponding example and reran the simulation using a new set of randomly selected parameters within the ranges specified in Tables 1, 2, and 3. Importantly, the numerical instabilities we refer to are not due to the use of the explicit Euler method, but instead stem from specific parameter combinations in the simulated chemical systems. For example, when the temperature is extremely low, the Arrhenius expression for the reaction rate can underflow, causing the reaction rate term to approach zero. This can result in degenerate system dynamics that are unrealistic or non-informative. The filtering process was therefore essential not only for removing unstable simulations but also for ensuring that the model is trained on physically meaningful and relevant trajectories. This contributes to the overall credibility and robustness of the foundation model development.
Remark 12.
The integration time step and the sampling period vary across CSTRs, BRs, and PFRs in this simulation study due to different kinetics and process dynamics. However, it is important to note that the “foundation model” can accommodate differing integration time steps and sampling periods for real-world reactors, provided the input dimensions remain consistent. This is due to the fact that the inputs to the neural network are merely numerical values, lacking any inherent physical interpretation for the network. Our approach does not require explicit architectural modifications because the RNN processes sequences of numerical values without relying on the physical interpretation of time steps. As long as the input dimension (number of features per timestep) and sequence length (maximum sequence length in a batch) remain consistent, the RNN can learn to map different reaction trajectories, regardless of the integration time step used during data generation. Specifically, The RNN is not inherently tied to a specific time discretization; it learns patterns based on numerical sequences rather than absolute time scales. Moreover, if different integration time steps result in varying trajectory lengths, padding or resampling would be necessary to ensure uniform input size. However, within a consistent input structure, the RNN can adapt to reactions evolving at different rates. However, while the RNN can implicitly learn representations across different time-step sizes, extreme differences in time scales (e.g., very stiff reactions) may still require additional considerations, such as adaptive sampling or time-step normalization.